OpenStudy (anonymous):

can someone please help me on this one Subtract the polynomials: (3x3 - 2x + 3) - (2x3 - 2)

OpenStudy (anonymous):

are you sure you types the first trinomial correctly?

OpenStudy (anonymous):

typed*

OpenStudy (anonymous):

yes, the 3x3 is 3x^3 do you get it or not... i can screet shot it

OpenStudy (neer2890):

just add the numbers of same power e.g., 3x^3-2x^3=x^3

OpenStudy (anonymous):

5x^3-2x+1 that's the answer

OpenStudy (anonymous):

THANKS @joanip_

OpenStudy (anonymous):

you're welcome

OpenStudy (neer2890):

@joanip_ i think you misunderstood the question it's \[3x^{3}-2x+3-(2x ^{3}-2)\] \[=3x ^{3}-2x+3-2x ^{3}+2\] \[=x ^{3}-2x+5\]

OpenStudy (neer2890):

i forgot to mention the like terms \[3x ^{3}-2x ^{3}-2x+3+2=x ^{3}-2x+5\]

OpenStudy (anonymous):

oh, i'm sorry :c

OpenStudy (neer2890):

it's okay.. :)